OpenStudy (anonymous):

if the supplement of an angle A is 5/2 of its complement, find the value of A

hartnn (hartnn):

u know the meaning of supplement and complement ?

OpenStudy (anonymous):

complement is 90-A and supplement is 180-A?? is that it??

hartnn (hartnn):

yup. so where are u stuck ?

OpenStudy (anonymous):

i don't know how to illustrate the equation..

hartnn (hartnn):

supplement = 180-A = (5/2) of complement = (5/2)* (90-A) 180-A=(5/2)*(90-A)<----equation can u solve that ?

OpenStudy (anonymous):

wait.. i'll try

OpenStudy (anonymous):

A=30 :)) thanks!!

hartnn (hartnn):

yup, 30 welcome :)